A heating element in a stove is designed to dissipate 2,542 W when connected to 240 V.

(a) Assuming the resistance is constant, calculate the current in the heating element if it is connected to 120 V.

(b) Calculate the power it dissipates at that voltage.
Answer #1

a) for the resistance of the heating element

as Power = V^2/R

R = 240^2/2542 = 22.66 Ohm

for the voltage, V = 120 V

current in the heating element = V/R

current in the heating element = 120/22.66

current in the heating element = 5.3 A

b) at 120 V

power dissipated = V^2/R

power dissipated = 120^2/22.66

power dissipated = 635.5 W